Liz drove the first half of a trip at 36 mi/h. At what speed should she cover the remaining half in order to average 45 mi/h for the whole trip
Answer by stanbon(75887)

Liz drove the first half of a trip at 36 mi/h. At what speed should she cover the remaining half in order to average 45 mi/h for the whole trip
-----
1st half DATA:
distance = x miles ; rate = 36 mph ; time = x/36 hrs.
----
Whole trip DATA:
distance = 2x miles ; rate = 45 mph ; time = 2x/45 hrs
--------------------
2nd half DATA:
distance = x miles;
time = (2x/45)-(x/36) = (72x-45x)/(45*36)
= 27x/(45*36) = x/(5*12) = x/60 hrs
rate = d/t = x/(x/60) = 60 mph
==================================
Ans: 60 mph
================
Cheers,
Stan H.
==========